# On $\mathcal{F}$-semi-transitive families of operators

### v6

@@ 4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}.$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\\ \{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that

### v7

# Disjoint supercyclicity in Banach algebras

@@ 1. Introduction and preliminaries
[18, Definition 4.1] Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that